13.29 m MTWA A 4.2.L volume of neon gas (Ne) is at a pressure of 4.1 atm and a temperature of 460 K. The atomic mass of neon is 20.2-/mol, and the ideal gas constant is R = 8.314J/mol-K The mass of the neon gas is closest to 9. a. 0.01717 kg b. 0.00964kg c 0.01636kg d. 0.01566 kg e. 0.009214.kg f0.01595.kg to search
Answer #1

Solution :-

From the ideal gas equation we have,

PV = nRT

where n is the number of moles of gas

n = PV/RT = (4.2×4.1)/(0.082×460) here R = 0.082 in J/litre-K

n = 0.46 moles

Also number of moles = mass/ atomic mass

Mass = atomic mass × number of moles = 20.2 × 0.46 = 0.009214 kg

Hence mass of neon gas is 8.09 grams